please help me here! janet and lynn live 8 blocks apart in the opposite directions from their office. If Lynn lives 1 block less than twice as far from the office as Janet does, how far does each girl live from the office?

Answer by Cintchr(481)

Janet and Lynn live 8 blocks apart ... J + L = 8
Lynn lives 1 block less than twice as far as Janet ... L = 2J - 1
J + L = 8 ... solve for J
- L = - L
J = 8 - L
Substitute this into L = 2J - 1
L = 2(8 - L) - 1
Distribute.
L = 16 - 2L - 1
Combine like terms
L = 15 - 2L
+2L = + 2L
3L = 15
divide by 3
L = 5 Lynn lives five blocks away.
J + L = 8
J + 5 = 8
- 5 = -5
J = 3 Janet lives three blocks away.